Session: May 2002 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Wed Jun 5 19:53:59 CEST 2002) Success
No problems.

Medicina (Poppi at Wed Jun 5 20:13:42 CEST 2002) Success
no problems

Urumqi (Chen Maozheng at Thu Jun 6 06:40:46 CEST 2002) Success

Yebes (Pablo de Vicente at Thu Jun 6 10:53:00 CEST 2002) Minor failures
1) There was a missuse of caltemps for S Band. The used value was 12.8 K whereas the correct value is 20.5 K. The error was due to changing the IF from IFB (normal position) to IFC (position for this experiment) for S band. Recalibration is needed with new figures. 2) Source not observed at 2002.155.23:47:47 3) Head 19 seems to be broken.

Effelsberg (A. Kraus at Thu Jun 6 14:03:55 CEST 2002) Minor failures
We had to stop twice because of a thunderstom and very strong rain: UT2058-2155 and UT1345-1600.

Noto (P.Cassaro at Thu Jun 6 16:17:02 CEST 2002) Success

Onsala (John Conway at Sat Jun 8 16:43:53 CEST 2002) Success
No known problems